OpenStudy (anonymous):

Generate the first 5 terms in the sequence using the explicit formula. y(n)=-5n-5 n in parentheses is subscript

OpenStudy (campbell_st):

so substitute n = 1 then 2, 3, 4, 5 n = 1 y(1) = -5(1) - 5 n = 2 y(2) = -5(2) - 5 just keep repeating till you get to n = 5

OpenStudy (anonymous):

Oh I get it. Thanks!